Find the slope of the tangent line at the incident point

3e to the power x = xy+xsquared+ysquared at (0,1)

Write an equation of the tangent line


1
Expert's answer
2021-05-19T08:30:50-0400
"(3e)^x=xy+x^2+y^2"

Differentiate both parts with respect to "x" . Use the Chain Rule


"\dfrac{d}{dx}((3e)^x )=\dfrac{d}{dx}(xy+x^2 +y^2)"

"(\ln3+1)(3e)^x=y+x\dfrac{dy}{dx}+2x+2y\dfrac{dy}{dx}"

Solve for "\dfrac{dy}{dx}"


"\dfrac{dy}{dx}=\dfrac{(\ln3+1)(3e)^x-2x-y}{x+2y}"



Find the slope of the tangent line at point "(0, 1)"


"slope=m=\dfrac{dy}{dx}|_{(0,1)}"

"=\dfrac{(\ln3+1)(3e)^0-2(0)-1}{0+2(1)}=\dfrac{\ln3}{2}"

The equation of the tangent line in point-slope form


"y-1=\dfrac{\ln3}{2}(x-0)"

The equation of the tangent line in slope-intercept form


"y=\dfrac{\ln3}{2}\cdot x+1"